Students pass a test if they score 50% or more. The marks of a large number of students were sampled and the mean and standard deviation were calculated as 42% and 8% respectively. Assuming this data is normally distributed, what percentage of student pass the test

1
Expert's answer
2022-02-14T08:33:32-0500

let X be the sore random variable.

"\begin{aligned}

&E(X)=42

\%=0.42 \\

&S

D(X)=8 \%=0.08 \\

&P(X>0.50)=?

\end{aligned}"

We know that if "X \sim N(\mu, \sigma)" then,

"\begin{aligned}

z&=\frac{X-\mu}{\sigma} \sim N(0,1) \\
&\text{So,}\ P(\frac{X-0.42}{0.08}>\frac{0.50-0.42}{0.08})\\

&=P(z>1)

\end{aligned}"

From z-table calculator.

"P(X>0.50)=P(z>1)=0.15866"

So "15.866 \%" students will pass the test.